a particleparticle moves along a straight line. a force acts on particle which produces a constant power it starts with initial velocity 3 and after moving a distance 252m its velocity is 6.find the time taken

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Power is constant hence,
a = (6-3)/t = 3/t m/s^2 (change in velocity)
From 2nd law of eqn. we get,
252 = 3*t + 0.5*(3/t)*t^2
252 = 3t + 1.5t
t = 252/4.5
t= 56 s